OpenStudy (anonymous):

find two positive real numbers whose product is a aximum and whose sum is 138

OpenStudy (anonymous):

a+b=138 ab ----> maximize

OpenStudy (anonymous):

no i am in precalc

OpenStudy (anonymous):

our options are a) 67,71 b) 69,69 c)74,64 d)78,60 e)57,81

OpenStudy (anonymous):

The highest point is the vertex, so that what we should find

OpenStudy (anonymous):

(138-b)b=138b-b^2 -b/2a

OpenStudy (dumbcow):

You could also use the fact that given a perimeter, the largest area is a square. So to maximize a*b, then a should equal b. a=b=69